OpenStudy (anonymous):

Nakim simplified 3 times the square root of 2x plus x times the square root of 8x minus 5 times the square root of 18x and got -10x times the square root of 2x for an answer. Part 1: Using complete sentences, explain what Nakim did wrong. Part 2: Show all your work to simplify the expression.

OpenStudy (anonymous):

I don't understand what to do to explain what he did wrong

OpenStudy (anonymous):

lets see if we can actually write it

OpenStudy (anonymous):

\[3\sqrt{2x}+x\sqrt{8x}-5\sqrt{18x}\] is that the original expression?

OpenStudy (anonymous):

Yes that is the original expression

OpenStudy (anonymous):

ok so lets see what we can do with it, and maybe we will see why the answer of \(-10\sqrt{2x}\) is wrong

OpenStudy (anonymous):

\(\sqrt{8x}=\sqrt{4\times 2x}=\sqrt{4}\times \sqrt{2x}=2\sqrt{2x}\) similarly \(\sqrt{18x}=3\sqrt{2x}\)

OpenStudy (anonymous):

therefore \[3\sqrt{2x}+x\sqrt{8x}-5\sqrt{18x}\] \[=3\sqrt{2x}+2x\sqrt{2x}-15\sqrt{2x}\]

OpenStudy (anonymous):

well Nakim is just an idiot, that is all first of all what parent names their kid "nakim"? so he/she is not responsible, probably just bad genes

OpenStudy (anonymous):

in any case if you combine like terms, you will get \[-12\sqrt{2x}+2x\sqrt{2x}\]

OpenStudy (anonymous):

because the like terms are \(3\sqrt{2x}\) and \(-15\sqrt{12}\) not the \(2x\sqrt{2x}\)

OpenStudy (anonymous):

haha I thought the same thing about the name! thank you very much!! you have help me a ton